In an isosceles right triangle , the square of the hypotenuse is 98cm^2.
Find the length of the congruent sides

Answer by ewatrrr(24785)

 
Hi,
Using the Pythagorean Theorem
x^2 + x^2 = 98cm^2
2x^2 = 98cm^22
x^2 = 49cm^2
x = 7cm